OpenStudy (anonymous):

A triangle has coordinates A (1, 5), B (-2, 1) and C (0, -4). What are the new coordinates if the triangle is rotated 90° clockwise around the origin?

OpenStudy (ivettef365):

for 90 clockwise it is (x,y) -> (y, -x)

OpenStudy (ybarrap):

$$ A~rotation~by ~90^\circ means:\\\\ \begin{bmatrix} A_x &A_y \\ B_x &B_y \\ C_x &C_y \end{bmatrix} \begin{bmatrix} 0 &-1 \\ 1 &\quad \!\! 0 \end{bmatrix} = \begin{bmatrix} A_y &-A_x \\ B_y &-B_x \\ C_y &-C_x \end{bmatrix} \\\\ \text{Therefore,} \\\\ \begin{bmatrix} \quad 1 &\quad 5 \\ -2 & \quad 1 \\ \quad 0 & -4 \end{bmatrix} \begin{bmatrix} 0 &-1 \\ 1 &\quad \!\!0 \end{bmatrix} = \begin{bmatrix} \quad 5 &-1 \\ \quad 1 &\!\!\quad 2 \\ -4 & \!\!\quad 0 \end{bmatrix} \\\\ Rotated~vectors~are~\vec{A}=<5,-1>\vec{B}=<1,2>\vec{C}=<-4,0> $$